OpenStudy (anonymous):

Solve this equation for A: A ÷ 2 = 4 A. 4 B. 2 C. 6 D. 8

OpenStudy (plainntall):

That equation can be expressed as a fraction A/2=4 what do you multiply both side by to get only A on one side?

OpenStudy (anonymous):

\[(\frac{ 2 }{ 1 })\frac{ A }{ 2 }=4(\frac{ 2 }{ 1 })\] so A=8 Letter D

OpenStudy (plainntall):

yes, exactly, good job